Which of the following equations have solution? justify your answers

a) x2 =3 (mod 137 ) this mean x to the 2nd power congruence to 3 mod 13 to the 7 power

b) x3 =4 (mod 115.239) this mean x to 3rd power congruence to 4 mod 11 to the 5 th power times 23 to the 9th power

c) x7=2(mod49) this mean x to the 7th power congruence to 2 mod 7 to the 2nd power

d) x7 = 18 (mod 49) this mean x to the 7th power congruence to 18 mod 7 to the 2nd power
Answer #1

a) 2 = 3 (mod 137) Consider x = 3 (mod 13). Write f(x)= 2²-3. So, we are looking for solutions to f(x) = 0 (mod 13). We have1 *54mod 119. 23°). Write f(x) = x - 4. Cost We examine f(x) = 0 (mod 115) and f(x) = 0 (mod 28?). Note that f(3) = 0 (mod 23c) x²= 2 (mod 49). By Fermats little theorem, x= x (mod 7) for any x. Consider f(x) = -2. Now f(x) = 0 (mod 7 ) a=2 (mod 7).d) 2 = 18 (mod 49). - Write f(x) = x1-18. Now, f(x)=0 (mody) () 2 = 4 (modr). Again, f (x) = Te and f (A) = 7= 0 (mod 7). So